OpenStudy (anonymous):

Teaneck High had an enrollment of 10900 in 1982 and an enrollment of 13000 in year 2000. What is the predicted enrollment of Teaneck High in 2012 if we assume a linear model? Predicted Enrollment

OpenStudy (anonymous):

Use those two points to find the equation of the line, then plug in the year you are looking for (2012). Does that help?

OpenStudy (anonymous):

well i did do that buy what number would be used for 2000 would it be 20

OpenStudy (anonymous):

Why not use 2000? Your two points can be (1982, 10900) and (2000, 13000).

OpenStudy (anonymous):

oh nvm then

OpenStudy (anonymous):

let me try it out then

OpenStudy (anonymous):

okay well i got the slope as 3/350

OpenStudy (anonymous):

so do i set it up as y-y1=m(x-x1)

OpenStudy (campbell_st):

you also need to use the initial population in your linear model.... the form should be y = mx + b m = slope, b = y-intercept slope m = ( change in population)/(change in time) then when you have the slope put it into y = mx + b use either point to find the initial population, (y-intercept) and then you'll have your linear relationship. Lastly substitute x = 30 to find the student population in 2012